# Free energy of the gas of spin 1/2 fermions beyond the second order and the Stoner phase transition

### v1

@@ 2. The formalism
valid for $x\gg 1$ (obtained by inverting the Sommerfeld expansion [28] of the function $f(\nu)$) is used, the formula (21) reproduces the textbook [4, 5] low temperature equilibrium condition (equivalent to the condition $\mu_{+}=\mu_{-}$) and leads to the well known prediction that the Stoner phase transition to the ordered state is continuous with divergent magnetic susceptibility characterized by the critical exponent $\gamma=1$ and a finite discontinuity of the heat capacity. (In fact, this continuous character of the transition is accidental: in the same approximation the transition is of first order in the system of spin $s>1/2$ fermions and/or if the space dimension is not 3.) If the correction $F^{(2)}$ is included, the transition becomes first order, at least at sufficiently low temperatures [27, 11, 13] in agreement with the arguments given in the past in [17]. However the computation of the complete order $(k_{\rm F}a_{0})^{3}$ correction to the ground state energy $E$ performed in [14, 15, 16] showed that the first order character of the transition at zero temperature, very clear in the second order approximation, is strongly weakened, at least as long as the contributions proportional to $k_{\rm F}r_{0}$ and $k_{\rm F}a_{1}$ (which, if the underlying interaction potential $V_{\rm pot}$ is “natural”, i.e. if all $a_{\ell}$, $r_{\ell}$, etc. are of the same order of magnitude, are of the same order, $(k_{\rm F}R)^{3}$, as the $(k_{\rm F}a_{0})^{3}$ correction) are not taken into account [16]. Below we extend the existing computations in two ways: we compute the complete (proportional to $(C_{0}^{\rm R})^{3}$, i.e. to $(k_{\rm F}a_{0})^{3}$) temperature dependent third order corrections to the free energy $F$ and, moreover, we show how to perform the ressumation of two infinite subsets of temperature dependent corrections to $F$ of which the first one is the finite temperature generalization of the subset of diagrams taken into account in (the last section of) ref. [18].
3. Order $(k_{\rm F}a_{0})^{2}$ and order $(k_{\rm F}a_{0})^{3}$ particle-particle corrections to $F$
We begin by recalling the computation of the order $(C_{0}^{\rm R})^{2}$ term $F^{(2)}$ performed in [29]. In agreement with the formulated prescription it is given by the single Feynman diagram shown in Figure 1. The corresponding analytical expression can be obtained by convoluting either two AA-”blocks” or two BB-”blocks” shown in the same Figure:
@@ 3. Order $(k_{\rm F}a_{0})^{2}$ and order $(k_{\rm F}a_{0})^{3}$ particle-particle corrections to $F$
To make the formulae resulting from convoluting $A$-blocks more transparent it will be convenient to introduce the following notation:
@@
The form of the distribution functions $n_{\pm}({\mathbf{p}})$ plays the role only in the second one of these two identities.
                $A(\omega_{l+1}^{B},\mathbf{q})=$                      $^{q-k,~l-n}$                      $_{k,~n}$                              $B(\omega_{l}^{B},\mathbf{q})=$                                    $^{k,~n+l}$                                    $_{k+q,~n}$
Figure 1: The order $C_{0}^{2}$ diagram contributing to the thermodynamic potential $F$ of the gas of spin $1/2$ fermions and two “elementary” one-loop diagrams (AA- and BB-”blocks”) out of which the second order and those higher order (in the $C_{0}$ coupling) contributions which are taken into account in this work are composed. Solid and dashed lines denote propagators of fermions with the spin projections ++ and -, respectively.
In the introduced notation the $A$-block (obtained in [29]) takes the form
@@
After the sum in (23) over the bosonic Matsubara frequencies $\omega^{B}_{l}=(\pi/\beta)l$ is performed using the standard formulae [9, 29] one gets
@@
The first term in (28) is divergent. The change of the variables ${\mathbf{q}}=2{\mathbf{s}}$, ${\mathbf{p}}_{1}={\mathbf{s}}-{\mathbf{t}}_{1}$, ${\mathbf{p}}_{2}={\mathbf{s}}-{\mathbf{t}}_{2}$ (the Jacobian equals 8) makes the innermost integral elementary and allows to write it in the form
@@
In [29] it has been found that it is convenient to evaluate (the finite part of) $F^{(2)}/V$ by substituting ${\mathbf{p}}_{1}={\mathbf{k}}_{1}$, ${\mathbf{q}}={\mathbf{k}}_{1}+{\mathbf{k}}_{2}$, ${\mathbf{p}}_{2}={\mathbf{p}}$ (the Jacobian is 1), replacing (by another change of the integration variable) $n_{-}({\mathbf{k}}_{1}+{\mathbf{k}}_{2}-{\mathbf{p}})$ with $n_{-}({\mathbf{p}})$ and then performing explicitly the integral over the cosine of the angle between $\mathbf{p}$} and ${\mathbf{k}}_{1}+{\mathbf{k}}_{2}$. This allows to represent the order $(k_{\rm F}a_{0})^{2}$ contribution to $F$ in the form
@@
The (dimensionless function) $L({\mathbf{k}}_{1},{\mathbf{k}}_{2})$ is given by the single integral
@@
Figure 2: The particle-particle and the particle-hole diagrams contributing in the order $C^{3}_{0}$ to the thermodynamic potential $F$.
If only the interaction term proportional to $C_{0}$ in (1) is taken into account, there are two Feynman diagrams contributing to the free energy $F$ in the third order. The first one, shown in the left panel of Figure 2, is termed the particle-particle ring diagram. Its contribution $F^{(3)pp}$ is given by the convolution of three $A$-blocks
@@
where “two other terms” means the terms in which the role of ${\mathbf{p}}_{1}$ is played by ${\mathbf{p}}_{2}$ and ${\mathbf{p}}_{3}$. It is good to make at this point a contact with the contribution of this third order particle-particle ring diagram to the ground state energy density $E/V$ obtained in [14, 15] (and in [16]) to which the expression (34) should reduce in the zero temperature limit, that is when the Fermi distribution functions are replaced by the theta functions. The contribution to $E/V$ of the particle-particle diagram was in [14, 15] given by two terms (Eq. (21) in [15]) whereas here it is given by the single term (three seemingly identical terms). The equivalence of the two approaches is ensured by the algebraic, i.e. independent of the precise forms of $N_{--}^{\mathbf{p}}$ and $\mathbf{p}$}] (recall that $\{{\mathbf{p}}\}=N_{--}^{\mathbf{p}}-N_{++}^{\mathbf{p}}$), identity which results from the symmetrization:
@@
After using the symmetry, i.e. taking only the content of the bracket and multiplying it by 3, it allows to rewrite the expression for $F^{(3)pp}/V$ as the sum of two terms which in the $T=0$ limit precisely reduce to the two terms, $G_{1}$ and $G_{2}$, which in [14, 15] contributed to $E/V$.
Naively, as all the three terms of (34) seem also identical, one is tempted to compute only one of them and multiply the result by three. $F^{(3)pp}/V$ would be in this way given by a single four-fold integral. This, as we have found, would lead to an incorrect result which in the zero temperature limit would not agree with the contribution of the particle-particle diagram to $E/V$ (this is precisely the problem which did no allow us to immediately extend the computation reported in [29]).
@@
One can now check that in the sum $F^{(1)}+F^{(2)}+F^{(3)pp}$ all the divergences (up to the order $(k_{\rm F}a_{0})^{3}$) cancel out (as will be seen, the contribution $F^{(3)ph}$ of the other diagram of Figure 2, which completes the order $(k_{\rm F}a_{0})^{3}$ contribution to $F^{(3)}$, is finite; this also follows from the computations of the order $(k_{\rm F}a_{0})^{3}$ corrections to $E/V$ performed in [14, 15]). Writing $\{{\mathbf{p}}_{i}\}=-1+\{{\mathbf{p}}_{i}\}^{\rm sub}$ in the first term in the square bracket in (37) allows to single out the divergent part of $F^{(3)}/V$. It is given by (to this order one can set in (37) $C_{0}=C_{0}^{\rm R}$; we also suppress the symbol $P$ of the Principal Value)
@@
in which $p_{\rm min}=|\Delta_{-}|$ and $p_{\rm max}=\Delta_{+}$ are determined by the condition that the zero of the argument of the Dirac delta treated as a function of the cosine of the angle between $\mathbf{p}$} and ${\mathbf{k}}_{1}+{\mathbf{k}}_{2}$ lies between 1-1 and +1+1. Since $\varepsilon_{\mathbf{p}}$ depends on $p^{2}$, the function $L_{\delta}({\mathbf{k}}_{1},{\mathbf{k}}_{2})$ can, of course, be obtained in a closed analytic form. The expression (39) is finite (as indicates its superscript) the ultraviolet convergence [^3] The singularity introduced by the factor $1/|{\mathbf{k}}_{1}+{\mathbf{k}}_{2}|^{2}$ is superficial because for ${\mathbf{k}}_{1}+{\mathbf{k}}_{2}={\mathbf{0}}$ vanishes also the logarithm in (31) while in (40) $p_{\rm min}=p_{\rm max}$. of the integrations being secured by the exponential suppression provided by the Fermi distribution functions $n_{\pm}$. We have also checked that the expression (39) evaluated for $T=0$ (so that the Fermi distribution functions can be replaced by the step functions) reproduces numerically in the entire range of polarizations $P$ the contribution to the ground state energy density of the third order particle-particle diagram of Figure 2 obtained in [14, 15].
@@ 4. Resummation of the contributions of the particle-particle diagrams
It turns out that the contribution to the free energy $F$ of the infinite series of Feynman diagrams composed of $N$-fold rings of the particle-particle $A$-blocks of Figure 1 can be summed in a closed form. Consider first the order $(k_{\rm F}a_{0})^{N}$ term of this series (the factor $(-1)^{N+1}$ is the same as in (6) - there are as many rearrangements of the $\hat{\psi}_{+}$ operators as of the $\hat{\psi}_{-}$ ones; the factor $1/N!$ in (6) is reduced to $1/N$ as there are (N1)!(N-1)! identical diagrams)
@@
in which, in order to regularize the integrals, the substitution ${\mathbf{p}}_{l}]\rightarrow[{\mathbf{p}}_{l}]+il\epsilon$ has been made. Using the Sochocki formula this can be then rewritten (assuming that $\epsilon>0$ - it will be seen that the result does not depend on the sign of $\epsilon$) in the form
@@
in which the integrals of the factors $\{{\mathbf{p}}_{j}\}/([{\mathbf{p}}_{n}]-[{\mathbf{p}}_{j}])$ are understood in the Principal Value sense. The experience with the order $(k_{\rm F}a_{0})^{2}$ and $(k_{\rm F}a_{0})^{3}$ contributions teaches that removing divergences amounts simply to replacing $\{{\mathbf{p}}_{j}\}$ by $\{{\mathbf{p}}_{j}\}^{\rm sub}$ in the first terms of the integrands of the integrals over ${\mathbf{p}}_{j}$-s (but not in the delta-terms) and $C_{0}^{N}$ in front by $(C_{0}^{\rm R})^{N}$. This would be obvious had the dimensional regularization been used to handle ultraviolet divergences - by definition it sets the integrals like $\int_{\mathbf{p}}({\rm const})$ to zero and, as is well known (see e.g. [7]), $C_{0}=C_{0}^{\rm R}$ to all orders, if such a regularization is used. After the change of the variables ${\mathbf{p}_{n}}={\mathbf{k}}_{1}$, ${\mathbf{q}}={\mathbf{k}}_{1}+{\mathbf{k}}_{2}$ the order $(k_{\rm F}a_{0})^{N}$ particle-particle diagram contribution to $F$ can be then neatly written in the form
@@
Refer to caption Refer to caption
Figure 3: The difference $(F(P)-F(0))^{pp}/N$ (in units $(3/5)\varepsilon_{\rm F}$) for $T=0$ and $T=0.2~\!T_{\rm F}\equiv 0.2~\!\varepsilon_{\rm F}/k_{\rm B}$ as a function of the polarization $P=(N_{+}-N_{-})/N$ for different values of the gas parameter $k_{\rm F}a_{0}$.
The zero temperature analog of this formula (i.e. representing the contribution of the particle-particle ring diagrams to the ground state energy density $E/V$) has been for $P=0$ first given by Kaiser [31] who in deriving it relied on combinatoric arguments. The formula which is the zero temperature analog of (47) for arbitrary polarization $P$ has been then written (by invoking the Kaiser’s reasoning) down in [18, 30] (see also [32]).
@@
The numerical procedure for evaluating the expression (30) described in details in [29] (the main trick is to construct - for given values of $t=k_{\rm B}T/\varepsilon_{\rm F}$ and $P$, which together, through (18), determine the chemical potentials - the interpolations of the functions of the parameter $\Delta$ into which the function (31) can be decomposed) can be used to evaluate also the expression (47). Figures 3 and 4 show for four different values of the temperature and several values of the gas parameter $k_{\rm F}a_{0}$ the difference $(F(P)-F(0))^{pp}/V$ (in units $(k^{3}_{\rm F}/3\pi^{2})(3/5)\varepsilon_{\rm F}$) obtained by adding to the zeroth order term (17) given explicitly by the formula (25) in [29] the contribution (47). In agreement with the result obtained in the last section of the work [18], one can observe that for $T=0$ the minimum of $(F(P)-F(0))^{pp}/V$ starts to move away from $P=0$ for $k_{\rm F}a_{0}=0.79$ indicating the continuous transition to the ordered state. As can be expected, with increasing temperature this critical value of the expansion parameter shifts towards larger values (0.85, 0.92 and 1.12 for $T=0.2~\!T_{\rm F}$, $0.3~\!T_{\rm F}$ and $0.5~\!T_{\rm F}$, respectively). One can also see that the minimum is back at $P=0$ for $k_{\rm F}a_{0}>0.96$ (at $T=0$) - this is the first order “reentrant” transition to the paramagnetic state observed in [20, 30] which is the consequence of the existence of the maximum of the energy density (at $T=0$) for $P=0$ treated as a function of $k_{\rm F}a_{0}$ shown by the dashed (blue) lines in the left panel of Figure 5. (From the physical point of view this reentrant transition is largely irrelevant as it occurs for the values of the gas parameter for which the formation of dimers prevails and the free energy of the metastable state is no longer physical.) This maximum of the energy density (indeed seen in experiments with cold gases [33]) which occurs close to the Feshbach resonance on its so-called BEC side (i.e. for large positive scattering length $a_{0}$) and the existence of which for higher temperatures (for which there is no phase transition) has been given a theoretical explaination (using a completely different approach) in [34] results here, as has been shown in [18, 20], from the appearance for $k_{\rm F}a_{0}>1.34$ of the simple pole in the “in-medium” particle-particle elastic scattering amplitude which can be interpreted as being due to the “in-medium” positive energy bound state of two fermions (of opposite spin projections). The dashed (blue) lines on the right panel of this figure and in Figures 6 and 7 illustrate how the contribution (47) to the free energy and its maximum change with the polarization and the temperature.
@@
As the right panel of Figure 3 and Figure 4 show, with increasing temperature the reentrant transition occurs for higher ($1.06$ and $1.46$ for $T=0.2~\!T_{\rm F}$ and $T=0.3~\!T_{\rm F}$, respectively and yet higher for $T=0.5~\!T_{\rm F}$) values of the expansion parameter. The maximal depth of the minimum (at which $P\neq 0$) of $F$ first slightly decreases with the growing temperature (up to $T\approx 0.2~\!T_{\rm F}$) and then increases with it. Similarly $P=1$ is for temperatures up to $T\approx 0.2~\!T_{\rm F}$ reached only for $k_{\rm F}a_{0}$ values approaching the one at which the reentrant transition takes place but for higher temperatures it is reached well before it.

### v2

@@ 2. The formalism
valid for $x\gg 1$ [obtained by inverting the Sommerfeld expansion [28] of the function $f(\nu)$] is used, the formula (21) reproduces the textbook [4, 5] low temperature equilibrium condition (equivalent to the condition $\mu_{+}=\mu_{-}$) and leads to the well known prediction that the Stoner phase transition to the ordered state is continuous with divergent magnetic susceptibility characterized by the critical exponent $\gamma=1$ and a finite discontinuity of the heat capacity. (In fact, this continuous character of the transition is accidental: in the same approximation the transition is of first order in the system of spin $s>1/2$ fermions and/or if the space dimension is different than 3.) If the correction $F^{(2)}$ is included, the transition becomes first order, at least at sufficiently low temperatures [27, 11, 13] in agreement with the arguments given in the past in [17]. However the computation of the complete order $(k_{\rm F}a_{0})^{3}$ correction to the ground state energy $E$ performed in [14, 15, 16] showed that the first order character of the transition at zero temperature, very clear in the second order approximation, is significantly weakened, at least as long as the contributions proportional to $k^{3}_{F}a^{2}_{0}r_{0}$ and $(k_{F}a_{1})^{3}$ (which, if the underlying interaction potential $V_{\rm pot}$ is “natural”, i.e. if all $a_{\ell}$, $r_{\ell}$, etc. are of the same order of magnitude, are of the same, $(k_{\rm F}R)^{3}$, order as the $(k_{\rm F}a_{0})^{3}$ correction) are not taken into account [16]. Below we extend the existing computations in two ways: we compute the complete (proportional to $(C_{0}^{\rm R})^{3}$, i.e. to $(k_{\rm F}a_{0})^{3}$) temperature dependent third order corrections to the free energy $F$ and, moreover, we show how to perform the resummation of two infinite subsets of temperature dependent corrections to $F$ of which the first one is the finite temperature generalization of the subset of diagrams taken into account in (the last section of) ref. [18].
3. The order $(k_{\rm F}a_{0})^{2}$ and the order $(k_{\rm F}a_{0})^{3}$ particle-particle corrections to $F$
We begin by recalling the computation of the order $(C_{0}^{\rm R})^{2}$ term $F^{(2)}$ performed in Ref. [29]. In agreement with the formulated prescription it is given by the single Feynman diagram shown in figure 1. The corresponding analytical expression can be obtained by convoluting either two $A$-blocks or two $B$-blocks shown in the same figure:
@@
To make the formulas resulting from convoluting $A$-blocks more transparent it will be convenient to introduce the following notation:
@@
The form of the distribution functions $n_{\pm}({\mathbf{p}})$ plays the role only in the second of these two identities.
Figure 1: The order $C_{0}^{2}$ diagram contributing to the thermodynamic potential $F$ (or to the ground-state energy $E$) of the gas of spin $1/2$ fermions and two “elementary” one-loop diagrams (AA- and $B$-blocks) out of which the second order and those higher order (in the $C_{0}$ coupling) contributions that are taken into account in this work are composed. Solid and dashed lines denote propagators of fermions with the spin projections ++ and -, respectively.
In the introduced notation the $A$-block (obtained in Ref. [29]) takes the form
@@
After the sum in (23) over the bosonic Matsubara frequencies $\omega^{B}_{l}=(\pi/\beta)l$ is performed using the standard formulas [6, 29] one gets
@@
The first term in (28) is divergent. The change of the variables ${\mathbf{q}}=2{\mathbf{s}}$, ${\mathbf{p}}_{1}={\mathbf{s}}-{\mathbf{t}}_{1}$, and ${\mathbf{p}}_{2}={\mathbf{s}}-{\mathbf{t}}_{2}$ (the Jacobian equals 8) makes the innermost integral elementary and allows to write it in the form
@@
In Ref. [29] it has been found that it is convenient to evaluate (the finite part of) $F^{(2)}/V$ by substituting ${\mathbf{p}}_{1}={\mathbf{k}}_{1}$, and ${\mathbf{q}}={\mathbf{k}}_{1}+{\mathbf{k}}_{2}$, ${\mathbf{p}}_{2}={\mathbf{p}}$ (the Jacobian is 1), replacing (by another change of the integration variable) $n_{-}({\mathbf{k}}_{1}+{\mathbf{k}}_{2}-{\mathbf{p}})$ with $n_{-}({\mathbf{p}})$ and then performing explicitly the integral over the cosine of the angle between $\mathbf{p}$} and ${\mathbf{k}}_{1}+{\mathbf{k}}_{2}$. This allows to represent the order $(k_{\rm F}a_{0})^{2}$ contribution to $F$ in the form
@@
The (dimensionless) function $L({\mathbf{k}}_{1},{\mathbf{k}}_{2})$ is given by the single integral
@@
Figure 2: The particle-particle and particle-hole diagrams contributing in the order $C^{3}_{0}$ to the thermodynamic potential $F$ (or to the ground-state energy $E$).
If only the interaction term proportional to $C_{0}$ in (1) is taken into account, there are two Feynman diagrams contributing to the free energy $F$ in the third order. The first one, shown in the left panel of Fig. 2, is termed the particle-particle ring diagram. Its contribution $F^{(3)pp}$ is given by the convolution of three $A$-blocks
@@
where “two other terms” means the terms in which the role of ${\mathbf{p}}_{1}$ is played by ${\mathbf{p}}_{2}$ and ${\mathbf{p}}_{3}$. It is good at this point to make a contact with the contribution of this third order particle-particle ring diagram to the ground state energy density $E/V$ obtained in Refs. [14, 15] (and in Refs. [16]) to which the expression (34) should reduce in the zero temperature limit, that is when the Fermi distribution functions are replaced by the theta functions. The contribution to $E/V$ of the particle-particle diagram was given in Refs. [14, 15] by two terms (Eq. (21) in Ref. [15]) whereas here it is given by the single term (three seemingly identical terms). The equivalence of the two approaches is ensured by the algebraic, i.e. independent of the precise forms of $N_{--}^{\mathbf{p}}$ and $\mathbf{p}$}] (recall that $\{{\mathbf{p}}\}=N_{--}^{\mathbf{p}}-N_{++}^{\mathbf{p}}$), identity which results from the symmetrization
@@
After using the symmetry, i.e. taking only the content of the bracket and multiplying it by 3, it allows to rewrite the expression for $F^{(3)pp}/V$ as the sum of two terms which in the $T=0$ limit precisely reduce to the two terms, $G_{1}$ and $G_{2}$, which in Refs. [14, 15] contributed to $E/V$.
Naively, as all three terms of (34) seem also identical, one is tempted to compute only one of them and multiply the result by three. $F^{(3)pp}/V$ would be in this way given by a single fourfold integral. This, as we have found, would lead to an incorrect result which in the zero temperature limit would not agree with the contribution of the particle-particle diagram to $E/V$ (this is precisely the problem that did not allow us to immediately extend the computation reported in Ref. [29]).
@@
One can now check that in the sum $F^{(1)}+F^{(2)}+F^{(3)pp}$ all the divergences [up to the order $(k_{\rm F}a_{0})^{3}$] cancel out (as will be seen, the contribution $F^{(3)ph}$ of the other diagram of Figure 2, which completes the order $(k_{\rm F}a_{0})^{3}$ contribution to $F^{(3)}$, is finite; this also follows from the computations of the order $(k_{\rm F}a_{0})^{3}$ corrections to $E/V$ performed in Refs. [14, 15]). Writing $\{{\mathbf{p}}_{i}\}=-1+\{{\mathbf{p}}_{i}\}^{\rm sub}$ in the first term in the square bracket in (37) allows to single out the divergent part of $F^{(3)}/V$. It is given by (to this order one can set in (37) $C_{0}=C_{0}^{\rm R}$; we also suppress the symbol $P$ of the principal value)
@@
in which $p_{\rm min}=|\Delta_{-}|$ and $p_{\rm max}=\Delta_{+}$ are determined by the condition that the zero of the argument of the Dirac delta treated as a function of the cosine of the angle between $\mathbf{p}$} and ${\mathbf{k}}_{1}+{\mathbf{k}}_{2}$ lies between 1-1 and +1+1. Since $\varepsilon_{\mathbf{p}}$ depends on $p^{2}$, the function $L_{\delta}({\mathbf{k}}_{1},{\mathbf{k}}_{2})$ can, of course, be obtained in a closed analytic form. The expression (39) is finite (as its superscript indicates) the ultraviolet convergence [^3] The singularity introduced by the factor $1/|{\mathbf{k}}_{1}+{\mathbf{k}}_{2}|^{2}$ is superficial because for ${\mathbf{k}}_{1}+{\mathbf{k}}_{2}={\mathbf{0}}$ vanishes also the logarithm in (31) while in (40) $p_{\rm min}=p_{\rm max}$. of the integrations being secured by the exponential suppression provided by the Fermi distribution functions $n_{+}({\mathbf{k}}_{1})$ and $n_{-}({\mathbf{k}}_{2})$. We have also checked that the expression (39) evaluated for $T=0$ (so that the Fermi distribution functions can be replaced by the step functions) reproduces numerically in the entire range of polarizations $P$ the contribution to the ground state energy density of the third order particle-particle diagram of Figure 2 obtained in Refs. [14, 15].
@@ 4. Resummation of the contributions of the particle-particle diagrams
It turns out that the contribution to the free energy $F$ of the infinite series of Feynman diagrams composed of $N$-fold rings of the particle-particle $A$-blocks of Figure 1 can be summed in a closed form. Consider first the order $(k_{\rm F}a_{0})^{N}$ term of this series [the factor $(-1)^{N+1}$ is the same as in (6) - there are as many rearrangements of the $\hat{\psi}_{+}$ operators as of the $\hat{\psi}_{-}$ ones; the factor $1/N!$ in (6) is reduced to $1/N$ as there are (N1)!(N-1)! identical diagrams]:
@@
in which, in order to regularize the integrals, the substitution ${\mathbf{p}}_{l}]\rightarrow[{\mathbf{p}}_{l}]+il\epsilon$ has been made. Using the Sochocki formula this can be then rewritten (assuming that $\epsilon>0$ – it will be seen that the result does not depend on the sign of $\epsilon$) in the form
@@
in which the integrals of the factors $\{{\mathbf{p}}_{j}\}/([{\mathbf{p}}_{n}]-[{\mathbf{p}}_{j}])$ are understood in the principal value sense. The experience with the order $(k_{\rm F}a_{0})^{2}$ and $(k_{\rm F}a_{0})^{3}$ contributions teaches that removing divergences amounts simply to replacing $\{{\mathbf{p}}_{j}\}$ by $\{{\mathbf{p}}_{j}\}^{\rm sub}$ in the first terms of the integrands of the integrals over ${\mathbf{p}}_{j}$-s (but not in the delta-terms) and $C_{0}^{N}$ in front by $(C_{0}^{\rm R})^{N}$. This would be obvious had the dimensional regularization been used to handle ultraviolet divergences –
Refer to caption Refer to caption
Figure 3: The difference $(F(P)-F(0))^{pp}/N$ (in units $(3/5)\varepsilon_{\rm F}$) for $T=0$ and $T=0.2~\!T_{\rm F}\equiv 0.2~\!\varepsilon_{\rm F}/k_{\rm B}$ as a function of the polarization $P=(N_{+}-N_{-})/N$ for different values (indicated in the panes) of the gas parameter $k_{\rm F}a_{0}$.
by definition it sets the integrals like $\int_{\mathbf{p}}({\rm const})$ to zero and, as is well known (see e.g. Ref. [8]), $C_{0}=C_{0}^{\rm R}$ to all orders, if such a regularization is used. After the change of the variables ${\mathbf{p}_{n}}={\mathbf{k}}_{1}$, and ${\mathbf{q}}={\mathbf{k}}_{1}+{\mathbf{k}}_{2}$ the order $(k_{\rm F}a_{0})^{N}$ particle-particle diagram contribution to $F$ can be neatly written in the form
@@
The zero temperature analog of this formula (i.e. representing the contribution of the particle-particle ring diagrams to the gas energy density $E/V$) has been for $P=0$ first given by Kaiser [31] who in deriving it relied on combinatoric arguments. The formula that is the zero temperature analog of (47) for arbitrary polarization $P$ has then been written (by invoking the Kaiser’s reasoning) down in Refs. [18, 30] (see also Ref. [32]).
@@
The numerical procedure for evaluating the expression (30) described in detail in Ref. [29] [the main trick is to construct – for given values of $t=k_{\rm B}T/\varepsilon_{\rm F}$ and $P$, which together, through (18), determine the chemical potentials – the interpolations of the functions of the parameter $\Delta$ into which the function (31) can be decomposed] can be used to evaluate also the expression (47). Figures 3 and 4 show for four different values of the temperature and several values of the gas parameter $k_{\rm F}a_{0}$ the difference $(F(P)-F(0))^{pp}/V$ [in units $(k^{3}_{\rm F}/3\pi^{2})(3/5)\varepsilon_{\rm F}$] obtained by adding to the zeroth order term (17) given explicitly by the formula (25) in Ref. [29] the contribution (47). In agreement with the result obtained in the last section of the work [18], one can observe that for $T=0$ the minimum of $(F(P)-F(0))^{pp}/V$ starts to move away from $P=0$ for $k_{\rm F}a_{0}=0.79$ indicating the continuous transition to the ordered state. As can be expected, with increasing temperature this critical value of the expansion parameter shifts toward larger values (0.85, 0.92 and 1.12 for $T=0.2~\!T_{\rm F}$, $0.3~\!T_{\rm F}$ and $0.5~\!T_{\rm F}$, respectively) of the gas parameter $k_{\rm F}a_{0}$. One can also see that the minimum is back at $P=0$ for $k_{\rm F}a_{0}>0.96$ (at $T=0$) – this is the first order “reentrant” transition to the paramagnetic state observed in Refs. [19, 30] which is the consequence of the existence of the maximum of the energy density (at $T=0$) for $P=0$ treated as a function of $k_{\rm F}a_{0}$ shown by the dashed (blue) lines in the left panel of Figure 5. (From the physical point of view this reentrant transition is largely irrelevant as it occurs for the values of the gas parameter for which the formation of dimers prevails and the free energy of the metastable state is no longer physical.) This maximum of the energy density (indeed seen in experiments with cold gases [33]), which occurs close to the Feshbach resonance on its so-called BEC side (i.e. for large positive scattering length $a_{0}$) and the existence of which for higher temperatures (for which there is no phase transition) has been given a theoretical explanation (using a completely different approach) in Ref. [34], results here, as has been shown in Refs. [18, 19], from the appearance for $k_{\rm F}a_{0}>1.34$ of the simple pole in the “in-medium” particle-particle elastic scattering amplitude which can be interpreted as being due to the existence of the “in-medium” positive energy bound state of two fermions (of opposite spin projections). The dashed (blue) lines on the right panel of this figure and in Figs. 6 and 7 illustrate how the contribution (47) to the free energy and its maximum change with the polarization and the temperature.
@@
As the right panel of Fig. 3 and Fig. 4 show, with increasing temperature the reentrant transition occurs for higher ($1.06$ and $1.46$ for $T=0.2~\!T_{\rm F}$ and $T=0.3~\!T_{\rm F}$, respectively and yet higher for $T=0.5~\!T_{\rm F}$) values of the expansion parameter. The maximal depth of the minimum (at which $P\neq 0$) of $F$ first slightly decreases with the rising temperature (up to $T\approx 0.2~\!T_{\rm F}$) and then increases with it. Similarly $P=1$ is for temperatures up to $T\approx 0.2~\!T_{\rm F}$ reached only for $k_{\rm F}a_{0}$ values approaching the one at which the reentrant transition takes place but for higher temperatures it is reached well before it.